Derive the least squares normal equations for the model :

\hat{y} = c_{1} x +c_{2}x^2
Answer #1

y = c_{1}x +c_{2}x^{2}

Using the principle of least squares, we have to determine the constants c1 and c2 so that

E =\sum \left ( y -c_{1}x -c_{2}x^{2} \right )^2 is minimum

Equating to zero, the partial derivates of E with respect to c1 and c2 as

\frac{\partial E}{\partial c_{1}} =0= -2\sum x \left ( y -c_{1}x -c_{2}x^{2} \right )

\frac{\partial E}{\partial c_{2}} =0= -2\sum x^2 \left ( y -c_{1}x -c_{2}x^{2} \right )

The normal equations are

\sum xy = c_{1}\sum x^2 +c_{2}\sum x^{3}

\sum x^2y = c_{1}\sum x^3 +c_{2}\sum x^{4}